OpenStudy (anonymous):

A company distributes free candies to all the students of x schools. Each school has (x + 1) classes. The number of students in each class is 3 more than the number of classes in each school. Each student is given 4 candies. Part A: Write an expression to show the total number of candies distributed by the company in x schools. (4 points) Part B: What would x(x + 1) represent? When simplified, what would be the degree and classification of this expression? (4 points) Part C: How can you calculate the total number of students in each school? (2 points)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x schools x+1 classes per school x+4 students per class how many students are there total?

OpenStudy (anonymous):

3+(x+1)

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

i'm not sure

jimthompson5910 (jim_thompson5910):

think of a numeric example, say we had 10 schools (x = 10) 11 classes per school (x+1 = 11) 14 students per class (x+4 = 14) how many students are there total?

OpenStudy (anonymous):

14

jimthompson5910 (jim_thompson5910):

that's the number of students in one class

jimthompson5910 (jim_thompson5910):

how many students are there total?

OpenStudy (anonymous):

x+ (x+1)+(x+4)

jimthompson5910 (jim_thompson5910):

you aren't adding them though

jimthompson5910 (jim_thompson5910):

go back to that numeric example and you'll see why addition doesn't work

OpenStudy (anonymous):

im really not sure

jimthompson5910 (jim_thompson5910):

say we had 3 classes and 10 students per class how many students are there total?

OpenStudy (anonymous):

30

jimthompson5910 (jim_thompson5910):

how did you compute that

OpenStudy (anonymous):

10*10*10 sorry

jimthompson5910 (jim_thompson5910):

no you were correct the first time, I just wanted to be sure you multiplied 3*10

OpenStudy (anonymous):

oh okay

jimthompson5910 (jim_thompson5910):

so say we had 10 schools (x = 10) 11 classes per school (x+1 = 11) 14 students per class (x+4 = 14) how many students are there total? Well in this specific numeric example, there are 10*11*14 = 1,540 students total.

jimthompson5910 (jim_thompson5910):

using that example, how do we answer part A?

OpenStudy (anonymous):

x(x+1)(x+4)

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

that's the number of students each student gets 4 candies, so we multiply that by 4 to get 4x(x+1)(x+4)

jimthompson5910 (jim_thompson5910):

which wraps up part A

OpenStudy (anonymous):

OKAY Im getting it

jimthompson5910 (jim_thompson5910):

that's good

OpenStudy (anonymous):

part b?

jimthompson5910 (jim_thompson5910):

What would x(x + 1) represent?

OpenStudy (anonymous):

the number of classes per school?

jimthompson5910 (jim_thompson5910):

that's x+1

jimthompson5910 (jim_thompson5910):

x is the number of schools

jimthompson5910 (jim_thompson5910):

so together, x*(x+1) means ???

OpenStudy (anonymous):

number of schools per class?

jimthompson5910 (jim_thompson5910):

say x = 10 schools, x+1 = 11 classes per school so x*(x+1) turns into 10*11

jimthompson5910 (jim_thompson5910):

10*11 represents what?

OpenStudy (anonymous):

i'm sorry i dont know

jimthompson5910 (jim_thompson5910):

10 schools 11 classes per school think of it like that example where we wanted to find the total number of students in 3 classes

OpenStudy (anonymous):

@jim_thompson5910 okay so when we mulplliy it right?

jimthompson5910 (jim_thompson5910):

giving you what

OpenStudy (anonymous):

this is not my answer, but it's like school is twice and classes *school?

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean

OpenStudy (anonymous):

is it schools per classes? ugh, im not sure

jimthompson5910 (jim_thompson5910):

think back to 3 classes, 10 students per class we got 30 students total ------------------------------------------------------- This part is just like that BUT we aren't finding the total number of students

jimthompson5910 (jim_thompson5910):

you're probably overthinking

OpenStudy (anonymous):

I know lol

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

okay so is it classes total?

jimthompson5910 (jim_thompson5910):

total number of classes, yes

OpenStudy (anonymous):

finally lol!

jimthompson5910 (jim_thompson5910):

hopefully you see why that is the case

OpenStudy (anonymous):

@jim_thompson5910 can you please help me? http://openstudy.com/study#/updates/53854472e4b03fd7b604666f

OpenStudy (anonymous):

Simplified it's x^2+x and the degree is 2 and a binomial... yes i see

OpenStudy (anonymous):

am i right?

jimthompson5910 (jim_thompson5910):

yep, you are correct

OpenStudy (anonymous):

we can calculate the total number of students in each school by finding x right?

jimthompson5910 (jim_thompson5910):

yeah if we knew x, we can calculate x(x+1)(x+4) to find the number of students total

OpenStudy (anonymous):

OKAY and thats THE ANSWER TO C

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks!!! so much !